# Colliding with Adversaries at ECML-PKDD 2025 Model Robustness Competition 1st Prize Solution

### v1

@@ 1 Introduction
Given that the RDSA (Random Distribution Shuffle Attack) method is used to generate adversarial examples for TopoDNN ([3],[7]), we opted for a method-aware robustness strategy. In particular, we applied an antiRDSA approach that perturbs data points using the same technique employed to generate the adversarial evaluation set, while preserving their original labels. As a result, our augmented training set includes multiple correctly labeled, perturbed adversarial examples.
@@ 3 Data Generation
The data generation methodology is largely based on the Random Distribution Shuffle Attack (RDSA) [6], which was employed by the competition committee to create adversarial examples using a TopoDNN model. As the original algorithm relies on TopoDNN predictions to generate adversaries, our modified version of RDSA (so called antiRDSA) excluded the model query step and produced several large data points to assist in training. This data points contain a significant number of perturbed inputs, each labeled with the correct class. Consequently, the resulting data consist of many examples that resemble adversarial instances but are annotated with the true label.
@@ 4 Robust Model
The architecture adopted for this robust model follows the structure of feature embeddings combined with a simple MLP (Multi-Layer Perceptron) for classification, consistent with the approach proposed in [1]. The feature embedding component performs a non-linear transformation of each of the 87 scalar input features into a representation vector. This process can be formally described using $f_{e}:\mathbb{R}\rightarrow\mathbb{R}^{8}$. An abstract picture of the overall architecture is shown in Figure 1.
@@
While state-of-the-art methods employ periodic, bin-based, or stochastic embedding mechanisms [1, 8, 2, 5], such approaches could not be applied due to constraints imposed by the competition format. Instead, we implemented a simpler two-layer ReLU-activated embedding, using separate weights for each of the three input feature types: $p_{T}$, $\phi$, and $\eta$. The embedding of each of the 87 features is defined as:
@@ 5 Discussion
Our model’s architecture builds on recent work showing the value of feature embedding for tabular data. We adapt the PLR-MLP [1] formulation, substituting its periodic PLR embedding layer with a ReLU-activated alternative due to format constraints in the competition. Rather than fully sharing embedding parameters across all features or assigning completely separate parameters to each, we take a middle path: we group input features by their three distinct data types and learn one shared set of embedding parameters per type. This design reduces model size and inference cost while retaining more representational flexibility than a single shared embedding, striking a balance between parameter efficiency and expressiveness.
For future work, we believe performance could be improved by approximately 0.02–0.03 units through the integration of more modern architectures, such as the original periodic embedding variants, transformer-based models [8, 2], and other advances in tabular data modeling. Those examples have demonstrated much better results than simple MLPs and similar models on numerous standard benchmarks of tabular datasets, including physics-related material like the Higgs and Higgs-small datasets. Additionally, exploring stochastic methodologies, including Bayesian variational deep learning techniques [4], may further enhance performance by increasing the model’s robustness to adversarial or distributional shifts.

### v2

@@ 1 Introduction
Given that the RDSA (Random Distribution Shuffle Attack) method is used to generate adversarial examples for TopoDNN ([kasieczka2019machine],[saala2025introduction]), we opted for a method-aware robustness strategy. In particular, we applied an antiRDSA approach that perturbs data points using the same technique employed to generate the adversarial evaluation set, while preserving their original labels. As a result, our augmented training set includes multiple correctly labeled, perturbed adversarial examples.
@@ 3 Data Generation
The data generation methodology is largely based on the Random Distribution Shuffle Attack (RDSA) [saala2025enforcing], which was employed by the competition committee to create adversarial examples using a TopoDNN model. As the original algorithm relies on TopoDNN predictions to generate adversaries, our modified version of RDSA (so called antiRDSA) excluded the model query step and produced several large data points to assist in training. This data points contain a significant number of perturbed inputs, each labeled with the correct class. Consequently, the resulting data consist of many examples that resemble adversarial instances but are annotated with the true label.
@@ 4 Robust Model
The architecture adopted for this robust model follows the structure of feature embeddings combined with a simple MLP (Multi-Layer Perceptron) for classification, consistent with the approach proposed in [gorishniy2022embeddings]. The feature embedding component performs a non-linear transformation of each of the 87 scalar input features into a representation vector. This process can be formally described using $f_{e}:\mathbb{R}\rightarrow\mathbb{R}^{8}$. An abstract picture of the overall architecture is shown in Figure 1.
@@
While state-of-the-art methods employ periodic
, bin-based, or stochastic embedding mechanisms [gorishniy2022embeddings, voskou2024transformers, gorishniy2021revisiting, rahat2024volvo], such approaches could not be applied due to constraints imposed by the competition format. Instead, we implemented a simpler two-layer ReLU-activated embedding, using separate weights for each of the three input feature types: $p_{T}$, $\phi$, and $\eta$. The embedding of each of the 87 features is defined as:
@@ 5 Discussion
Our model’s architecture builds on recent work showing the value of feature embedding for tabular data. We adapt the PLR-MLP [gorishniy2022embeddings] formulation, substituting its periodic PLR embedding layer with a ReLU-activated alternative due to format constraints in the competition. Rather than fully sharing embedding parameters across all features or assigning completely separate parameters to each, we take a middle path: we group input features by their three distinct data types and learn one shared set of embedding parameters per type. This design reduces model size and inference cost while retaining more representational flexibility than a single shared embedding, striking a balance between parameter efficiency and expressiveness.
For future work, we believe performance could be improved by approximately 0.02–0.03 units through the integration of more modern architectures, such as the original periodic embedding variants, transformer-based models [voskou2024transformers, gorishniy2021revisiting], and other advances in tabular data modeling. Those examples have demonstrated much better results than simple MLPs and similar models on numerous standard benchmarks of tabular datasets, including physics-related material like the Higgs and Higgs-small datasets. Additionally, exploring stochastic methodologies, including Bayesian variational deep learning techniques [panousis2021wta], may further enhance performance by increasing the model’s robustness to adversarial or distributional shifts.